### v2

# Forge: Foundational Optimization Representations from Graph Embeddings

Abstract
Combinatorial optimization problems are ubiquitous in science and engineering, yet learning-based approaches to accelerate their solution often require solving a large number of hard-to-solve optimization instances to collect training data, incurring significant computational overhead. Existing methods require training dedicated models for each problem distribution for each downstream task, severely limiting their scalability and generalization. In this work, we introduce Forge, a method of pre-training a vector-quantized graph autoencoder on a large and diverse collection of mixed-integer programming (MIP) instances in an unsupervised fashion without dependency on their solution. The vector quantization process creates discrete code assignments that act as a vocabulary to represent optimization instances. We evaluate our approach under both supervised and unsupervised settings. For the unsupervised setting, we demonstrate that Forge embeddings effectively differentiate and cluster unseen instances. For the supervised setting, we fine-tune Forge embeddings and show that a single model predicts both the variables for warm-starts and integrality gaps for cut-generation across multiple problem type distributions. Both predictions help improve performance of a state-of-the-art, commercial optimization solver. Finally, we release our code and pre-trained Forge weights to encourage further research and practical use of instance-level MIP embeddings at https://github.com/skadio/forge/.
1 Introduction
Combinatorial Optimization (CO) problems are fundamental in science and engineering with widespread applications across multiple domains, including logistics, energy systems, network design, and recommendations (Gabriel Crainic et al., 2021; Miller et al., 2023; Kadıoğlu et al., 2024). Traditionally, CO problems have been addressed using carefully designed meta-heuristics and sophisticated optimization solvers. These classical approaches often demand significant domain expertise and computational resources, especially when problem size and complexity scales up.
Recent advances in Machine Learning (ML) have introduced promising alternatives for solving CO problems. The approaches for ML-guided optimization fall mainly under two categories: 1) end-to-end models that attempt to predict solutions (or objective) without depending on meta-heuristics or solvers, and 2) hybrid methods that 2a) replace computationally intensive components of traditional optimization solvers with learned models, e.g., learning to predict strong branching heuristics (Gasse et al., 2019), and 2b) guide meta-heuristics with learning from feedback (Cai et al., 2025b). For a comprehensive overview of ML-guided CO, we refer readers to the survey by Bengio et al. (2021).
Despite their potential, many of these learning-based methods face practical limitations. A significant drawback of these learning approaches is their heavy dependency on offline training. Training is computationally costly, depends on carefully curating training datasets with desired properties and distributions of the underlying CO instances, and has limited generalization. Moreover, training often depends on using optimization solvers in the first place to create labeled datasets, which defeats the purpose of improving solving for very hard instances that optimization solvers cannot deal with today. Adapting learning-based methods to new distributions and domains remains a challenge, and hence, designing foundational optimization representations in an unsupervised fashion applicable to multiple optimization tasks are a much-needed alternative. This is what we study in this paper.
Our main motivation is grounded in the exceedingly successful ML methods in other modalities such as foundational text and image embeddings. This raises a natural question for optimization: can we leverage the large number of publicly available mixed-integer programming (MIP) instances to develop a pre-trained, general-purpose, foundational model for MIP representation that is useful across multiple optimization tasks across varying sizes and problem domains?
The success of current ML-based approaches for optimization problems makes this direction imminent. For example, Zhou et al. (2023) proposes a meta-learning framework capable of generalizing across variants of vehicle routing problems of different sizes, albeit it is limited to one specific optimization problem, i.e., routing. Similarly, Cai et al. (2025a) introduces a multi-task framework based on a shared model for backdoor prediction and solver configuration, however, the shared model is trained separately per problem. Likewise, Li et al. (2025) proposes an LLM-based evolutionary framework to generate a large set of diverse MIP problems, but remains supervised, and requires a large number of apriori solved instances. While these are promising, the existing work either generalizes across multiple tasks but remains specific to one problem type, or scales to different sizes and problem variants but is specific to one task, or remains supervised. We envision a foundational model to create MIP embeddings that are generalizable to multiple tasks, across different problems, trained in an unsupervised fashion, without dependency on solving hard optimization problems.
When we consider Natural Language Processing (NLP) and Computer Vision (CV), foundational models have emerged through unsupervised or self-supervised training due to the abundance of data. CO problems, while also having a large amount of data, span highly heterogeneous problem types (e.g., Set Cover versus Combinatorial Auction) and present significant variability within each problem type. Most ML-based approaches to CO rely on Graph Neural Networks (GNN), as proposed in (Gasse et al., 2019), which are effective at capturing local variable- and constraint-level information but struggle to capture meaningful global information at the instance level, as studied in detail by Feng et al. (2025) where it was shown that GNNs often struggle to capture the global structure necessary for solving CO problems due to their inherent locality bias. As such, while embeddings for several other modalities such as words, images, and audio exist and have become widespread, to date, there exists no general-purpose MIP embeddings at the instance-level.
With this vision in mind, we propose Forge: Foundational Optimization Representations from Graph Embeddings. Forge is a foundational model to generate useful MIP embeddings through a pre-training framework that learns structural representations at the instance level from a broad distribution of MIP instances without requiring access to their solutions.
To achieve this, we incorporate two key ideas; one originating from NLP and the other from CV. As in NLP, we seek to build a vocabulary to represent the latent space of optimization problems, which enables instance-level representations. As in CV, we leverage vector quantization to preserve global information, which overcomes the issue of GNNs used for CO in previous works. By extending these two crucial insights into the CO context, we make the following concrete contributions:
  • We propose Forge, a general-purpose foundational model to generate MIP embeddings. We show that Forge effectively captures both local and global structures critical to optimization. Unlike previous work, with a single pre-trained Forge model, we provide embeddings at various levels; both instance-level representations, i.e., a single vector representation per MIP instance, and fine-grained variable- and constraint-embeddings.
  • For the unsupervised setting, we show that our instance level MIP embeddings clusters previously unseen instances across multiple problem types with high accuracy.
  • For the supervised setting, we show that pre-trained Forge embeddings can be fine-tuned for various downstream tasks using a small number of labeled data. We evaluate Forge on two radically different tasks to showcase its versatility: predicting variables for warm-starts and predicting integrality gap for cut-generation. Notably, a single pre-trained Forge model is fine-tuned and applied across diverse problem types including Combinatorial Auctions (CA), Set Covering (SC), Generalized Independent Set (GIS), and Minimum Vertex Cover (MVC).
  • To enhance traditional optimization solvers, we integrate predictions from Forge into Gurobi (Gurobi Optimization, LLC, 2024), a state-of-the-art commercial solver, and demonstrate consistently lower primal gaps across a range of problem domains and sizes.
  • To enhance ML-guided optimization methods, we incorporate Forge embeddings into PS-Gurobi (Han et al., 2023), enhancing its performance.
The remainder of the paper is structured as follows: we start with a brief background of CO in Section §2 followed by the Forge methodology in Section §3. Sections §4 and §5 show experiments to demonstrate the efficacy of Forge in unsupervised and supervised settings respectively.
Figure 1: Forge: The overall unsupervised approach for learning MIP embeddings. Forge uses a vector quantized graph autoencoder (C-D) to reconstruct node features and edges and is pre-trained across a diverse set of problem types and difficulty levels, enabling it to learn generalizable structural representations without requiring access to optimal solutions.
2 Background: Mixed-Integer Programming
We deal with solving Mixed-Integer Programming (MIP) instances that formulates combinatorial optimization problems of the form:
$f(x)=\min\{c^{T}x\mid Ax\leq b,x\in\mathbb{R}^{n},x_{j}\in\mathbb{Z}\ \forall j\in I\}$
where $f(x)$ is the objective function, and $A\in\mathbb{R}^{m\times n},b\in\mathbb{R}^{m},c\in\mathbb{R}^{n}$, and the non-empty set $I\subseteq{1,...,n}$ indexes the integer variables.
The Linear Programming (LP) relaxation of a MIP $x_{lp}$ is obtained by relaxing integer variables to continuous variables, i.e., by replacing the integer constraint $x_{j}\in\mathbb{Z}\ \forall j\in I$ to $x_{j}\in\mathbb{R}\ \forall j\in I$. The LP relaxation is an essential part of the branch-and-bound algorithm for providing bounds.
The integrality gap measures how much worse the optimal solution of the LP relaxation when compared to the optimal solution of the original MIP.
3 Forge: Unsupervised Approach to Learn MIP Embeddings
MIP instances are typically represented as a bipartite graph between variables and constraints augmented with node features (Gasse et al., 2019; Ferber et al., 2022; Yau et al., 2024; Chen et al., 2023). This is then followed by training a GNN in a supervised fashion for a specific downstream task on a certain problem class. For example, in Han et al. (2023), the GNN is used for predicting variables for warm-starts trained on Set Cover (SC) and Independent Set (IS) problems. Similarly, in Li et al. (2025), a GNN is used for predicting the integrality gap. Numerous variants follow this template (see §6). Notice that, all of these methods require supervision and do not yield a general-purpose MIP embeddings at the instance level. Taking this a step further; our goal is to learn the structure of MIP instances in an unsupervised manner.
Figure 1 presents our overall architecture, which is composed of these main building blocks:
A) MIP-to-BP: Given a MIP instance, we start with its bipartite (BP) representation and node features. Each node in this bipartite graph represents a constraint or a variable, with edges indicating which variables are part of which constraints. Each node is also associated with node features and each edge is weighted by the coefficient of the variable in the constraint. Node features are typically extracted from the internal branch-and-bound search tree when solving an instance. For example, there are 18 input node features used in Gasse et al. (2019). We do not attempt to solve or depend on the solution of the instance. Instead, Forge only uses the basic properties of the input instance. For each constraint node, we introduce 4 features composed of its sense (i.e., >>, << or ==) and the RHS value. For each variable node, we introduce 6 features composed of its type (integer, binary, continuous), upper/lower bound, and the coefficient in the objective function. In total, for each node, we obtain a vector of size 10, padded with zeros accordingly based on node type (Figure 1(A)).
B) BP-to-GNN: This bipartite graph with 10-dimensional input node features is then passed into a GNN, similar to previous works, to generate embeddings for each constraint and variable node. More specifically, Forge uses two GraphSage (Hamilton et al., 2017) layers that project each input node into a $d$ dimensional embedding space.
As discussed earlier, while GNNs are good at capturing local variable- and constraint-level information, they struggle to capture meaningful global information at the instance level due to their inherent locality bias (Feng et al., 2025). Preserving global structure is important in CO problems, especially to generalize across problem types (Figure 1(B)).
C) Vector Quantized Codebook: To preserve global structure, we introduce a vector quantized codebook with $k$ discrete codes. These codes act as a “vocabulary”, akin to language models, across MIP instances of various domains and difficulties, thereby preserving global structure. The design follows the approaches developed in computer vision (Van Den Oord et al., 2017; Yu et al., 2022; Lee et al., 2022) and the structure-aware graph tokenizer extension in Yang et al. (2024) (Figure 1(C)).
D) GNN-to-CW: GNN embeddings are passed into a vector quantizer which consists of a codebook with $k$ codes. The codebook maps each variable and constraint node to a discrete code. Each code is then mapped into a $d$ dimensional codeword (CW), yielding constraint and variable CW representations. These representations are the same dimension as the hidden GNN layers (Figure 1(D)).
E) CW-to-BP: We use the CW corresponding to each constraint and variable node to reconstruct the original bipartite representation of the MIP instance. These codewords are passed into a linear node feature decoder and a linear edge decoder to reconstruct the input bipartite graph. By doing so, we obtain an unsupervised method that learns from the structure of MIP instances (Figure 1(E)).
F) Loss Function: The loss function minimizes the edge reconstruction loss, the node feature reconstruction loss and losses related to the vector quantization. Concretely, the loss function is:
$\mathcal{L}=\mathcal{L}_{Rec}+\mathcal{L}_{Codebook}+\mathcal{L}_{Commitment}$ (1)
where given $N$ nodes, input node feature $v_{i}\ \forall i\in N$, the adjacency matrix $A$ and a matrix $\hat{X}$ composed of reconstructed input features $v_{i}$, the reconstruction loss, $\mathcal{L}_{Rec}$, the codebook loss, $\mathcal{L}_{Codebook}$, and commitment loss, $\mathcal{L}_{Commitment}$ are given by:
$\mathcal{L}_{Rec}=(A-\hat{X}\hat{X^{T}})^{2}+\frac{1}{N}\sum_{i=1}^{N}(\hat{v_{i}}-v_{i})^{2}$ (2)
$\mathcal{L}_{Codebook}=\frac{1}{N}\sum_{i=1}^{N}\|sg[h_{i}]-cw_{i}\|_{2}^{2}$ (3)
$\mathcal{L}_{Commitment}=\frac{\alpha}{N}\sum_{i=1}^{N}\|sg[cw_{i}]-h_{i}\|_{2}^{2}$ (4)
Here, $sg[.]$ is the stop-gradient operator, $h_{i}$ is the hidden layer representation of node $i$ after the GNN forward pass and $cw_{i}$ is the codeword corresponding to the code that node $i$ has been assigned.
Intuitively, the codebook loss in Eq. 3 can be interpreted as k-means clustering, where the codewords $cw_{i}$ (akin to cluster centroids) are moved closer to the node embeddings $h_{i}$ and the node embeddings $h_{i}$ are fixed in place due to the stop-gradient operator. Conversely, the commitment loss in Eq. 4 fixes the codewords $cw_{i}$ using the stop-gradient operator, and instead, moves the embeddings $h_{i}$ towards the codewords. The hyperparameter $\alpha$ weighs the importance of the commitment loss.
Figure 2: (A) Each node in the bipartite graph representation of the MIP instance is assigned a discrete code. (B) The distribution of these assigned codes yields the embedding of the MIP instance. The embedding dimension is the size of the codebook, i.e., $vocabulary=\lvert codebook\rvert$. This example uses 5 codes leading to the MIP embedding $\vec{emb}=[3,2,3,2,0]$.
Once Forge is trained in this unsupervised manner across a corpus of MIP instances, we obtain:
Local representations, where each node in the bipartite graph is assigned a discrete code. Each code is further mapped to a codeword which becomes the variable and/or constraint embedding.
Global representations, where we leverage the distribution of codes for the given MIP instance. As shown in Figure 2, each MIP instance is represented with an embedding of size $vocabulary=\lvert codebook\rvert$ where each value denotes the frequency of codes present in the instance. Figure 2 shows a codebook of size 5 (for brevity) and the resulting MIP embedding $\vec{emb}=[3,2,3,2,0]$.
Note that this form of unsupervised pre-training aims at learning the structure of MIP instances and is different from methods which aim to cast the discrete objective function of a MIP instance as a differentiable one and learn the solution using gradient descent in an end to end manner (Sanokowski et al., 2024; Karalias & Loukas, 2020; Bu et al., 2024).
Next, we start with an initial investigation of the effectiveness of Forge embeddings when clustering unseen instances across various problem domains (§4), and then explore how they can enhance solving MIP instances via warm-starts and cut-generation(§5).
4 Initial Analysis of Forge Embeddings
Given the absence of methods in the literature that can provide general-purpose MIP embeddings at the instance level, with the exception of earlier works that depend on creating hand-crafted features to classify MIP instances (e.g., (Kadioglu et al., 2010)), we start our initial analysis with a comparison of Forge embeddings against two (ablation) baselines when clustering unseen instances.
Our goal is to investigate both the accuracy of the clustering (quantitative) as well as visual inspection (qualitative). For this purpose, we train Forge on a set of MIP instances from MIPLIB (Gleixner et al., 2021), and test it on unseen instances from Distributional MIPLIB (D-MIPLIB) (Huang et al., 2024). While MIPLIB is a mixed dataset, D-MIPLIB is categorized into different problem classes. This serves as ground truth label when we treat each problem class as a cluster to evaluate our embeddings.
Figure 3: 2-D PaCMAP visualization of D-MIPLIB instances using MIP embeddings from (a) the mean readout of node embeddings from the GNN in Forge, (b) the two-hop label propagation of input node features, and (c) the distribution of discrete codes produced by Forge. Normalized Mutual Information (NMI), averaged over 10 runs of $k$-means clustering, is reported for each method.
Training: We use 600 instances from MIPLIB, sorted by size to ensure the resulting bipartite graphs fit on GPU memory. We further generate two more instances from each MIPLIB instance by randomly dropping 5% and 10% of constraints. Note that dropping constraints does not impact feasibility of the instance. In total, we obtain 1,800 MIP instances to train Forge in an unsupervised fashion as described earlier (§3). We use two GraphSage layers with 1024 dimensions for the hidden GNN layers ($d$) and a codebookcodebook with 5000 codes (i.e., the size of the vocabulary).
Testing: We evaluate Forge on 1,050 instances from D-MIPLIB organized into 21 problem type-difficulty pairs. These are distributed as Set Cover (easy, medium, hard), Maximum Independent Set (easy, medium), Minimum Vertex Cover (easy, medium, hard), Generalized Independent Set (easy, medium, hard, very-hard, very-hard2, ext-hard), Combinatorial Auction (very-easy, easy, medium, very-hard, very-hard2), Item Placement (very-hard) and Maritime Inventory Routing Problem (medium), thereby covering a broad spectrum of problem types, sizes and difficulty levels.
MIP Embeddings: These 1,050 instances are passed through Forge to generate one embedding vector per MIP instance using the distribution of discrete codes as shown in Figure 2. For comparison, we consider two other alternatives. As a baseline approach, we use the Mean Readout utilizing the GNN embeddings of the trained Forge model. The Mean Readout method generates the MIP embedding of an instance by averaging all node features from the GNN hidden layer representation from Figure 1(B). This baseline also behaves as an ablation for Forge without vector quantization. Given the weakness of GNN embeddings in capturing global structure (Feng et al., 2025), we expect this method to perform poorly. Alternatively, starting from the 10-dimensional static node features of the bipartite graph (Fig. 1(A)), we perform two-hop label propagation (Zhu & Ghahramani, 2002) and average the resulting node vectors. Note that global, static instance descriptors (e.g., (Kadioglu et al., 2010)) cannot be used since within D-MIPLIB, every instance in a category shares identical coarse statistics (number of variables, constraints, etc.), rendering them non-informative.
Clustering Visualization: Figure 3 visualizes these embeddings vectors, projected into two dimensions using PaCMAP (Wang et al., 2021). Each dot represents an instance colored by its problem type. As expected, the Mean Readout method looses the global structure leading to an arbitrary clustering. The Label Propagation performs much better capturing some structure, while some problem domains and sizes are mixed in the clustering. Forge embeddings performs the best where each problem type, such as Set Cover, Combinatorial Auction etc., are separated cleanly. Moreover, within each problem domain, the difference in problem difficulty is also observed from the easy to medium and hard categories. Recall again that our Forge model is not trained on these instances. We are able to train on MIPLIB instances and capture the structure of unseen D-MIPLIB instances.
Clustering Accuracy: For quantitative evaluation, we run k-means with 21 clusters, expecting one cluster for each problem type. We calculate the normalized mutual information score (NMI) between then ground truth problem type assignment and the k-means clusters generated by each method. The NMI scores are averaged over 10 runs of k-means clustering. If the predicted clusters are identical to the original problem types, then NMI score would be 1.0, and 0.0 otherwise.
As qualitatively observed from the 2D visualization, the Mean Readout performs poorly with an NMI score of 0.087. This is only slightly better than random cluster predictions with an NMI score of $1/21=0.047$. This is likely due to over-smoothing due to averaging dense GNN embedding vectors across a large number of constraints and variables. In contrast, the Label Propagation performs better with an NMI score of 0.790. This method operates directly on the sparse input features of the bipartite graph, thereby avoiding the over-smoothing problem encountered in deep GNN layers. The best performance is achieved by Forge with an NMI score of 0.843. Interestingly, the mean readout embeddings, which perform very poorly, are generated from the same trained GNN within Forge. By utilizing the distribution of the discrete codes of constraints and variables, Forge circumvents the over-smoothing issue and captures the global structure of unseen MIP instances.
5 Experiments
Having demonstrated that Forge embeddings reliably cluster unseen instances from diverse problem classes, we now shift to supervised evaluations more directly aimed at improving MIP solving. Accordingly, we design the next suite of experiments on:
  • Downstream tasks that are commonly applied in the literature to enable us with a fair comparison and provide utility to enhance MIP solving.
  • Downstream tasks that are radically different from each other, where we use the same Forge model, to validate its general applicability across tasks, problems, and sizes.
  • Downstream tasks that are agnostic to the underlying MIP solver, i.e., we do not depend on internal access to specific solver procedures (e.g., within the branch-and-bound tree).
We therefore study two different downstream tasks: predicting the integrality gap of a given MIP instance, as also used in (Li et al., 2025), and predicting variables for warm-starts, as also used in (Han et al., 2023), in Sections § 5.2 and § 5.3 respectively. Both tasks serve as primal heuristics to enhance MIP solving by obtaining better solutions faster. The integrality gap is used to generate a pseudo-cut added to the original problem formulation to tighten its bound. The warm-start variables are used to provide hints to the solver to start the solving process from a promising initial assignment.
A critical aspect of these experiments is that we use the same pre-trained Forge model to obtain general-purpose MIP embeddings that are then fine-tuned on a small number of labeled data to learn prediction heads for completely different tasks (as shown in Figure 4(A)). An analogy from NLP would be to generate pre-trained word embeddings using a foundational model, and then fine-tune prediction heads for entity extraction in a specific domain, e.g., finance, using a small set of labeled data. Our goal is to replicate the success of foundational models in NLP and CV in the context CO.
We begin with an overview of our training setup to create our pre-trained Forge model and then discuss its applications to predicting the integrality gap for pseudo-cut generation and predicting warm-start variables for guiding the search.
5.1 Training Setup for Foundational Forge Model
The setup for training Forge is identical to the setting in the initial clustering analysis, with two GraphSage layers with $d=1,024$ dimensions and $k=5000$ codes. The only difference is that we pre-train on the structure of the 1,800 MIPLIB instances as well as the 1,050 D-MIPLIB instances. In total, Forge is trained on a corpus of 2,850 MIP instances.
For these experiments, we train on the ml.g5.xlarge AWS instance with a GPU with 24 GB of memory. Inference experiments were run on the ml.c5.12xlarge instance with 48 cores and 96 GB of RAM. To ensure consistency and fairness, all experiments were executed with Gurobi, a state-of-the-art commercial MIP solver (Gurobi Optimization, LLC, 2024), restricted to a single thread and a time limit of 3600 seconds. Unsupervised pre-training and integrality gap fine-tuning are run for 10 epochs with a learning rate of $10^{-4}$, while the warm-start prediction task is trained for 25 epochs using a learning rate of $10^{-5}$.
5.2 Integrality Gap Prediction & Pseudo-Cut Generation
Integrality Gap Prediction: As covered earlier (§ 2), the integrality gap measures the ratio between the optimal solution of the LP relaxation and the optimal solution of the original integer program. Intuitively, this gap quantifies the quality of the approximation offered by the LP relaxation. A smaller gap means the LP relaxation is a good approximation and is close to the value of the best integral solution. Here, we are interested in predicting the integrality gap of a given MIP instance, without solving for its optimal integral solution value, as also studied in (Li et al., 2025).
Pseudo-Cut Generation: If we can predict the integrality gap, then we can generate a pseudo-cut and add it as an additional constraint to the model to immediately bound the optimal objective value from the initial LP relaxation. Note that, this cut is not guaranteed to be a valid cut at all times, hence it is a pseudo-cut. If the integrality gap prediction is incorrect, it risks over (or under) estimating the best objective value. This makes integrality gap prediction a challenging problem.
Figure 4: (A) The supervised fine tuning flow for Forge. Integrality gap predictions are generated by using a mean readout and warm start probabilities are generated for each variable using a sigmoid function. (B) Distribution of integrality gaps for different problem type-difficulty pairs. (C) Procedure for constructing triplets for the triplet loss for warm start prediction. Each MIP instance is solved to obtain the top 5 solutions. Nodes appearing in the same number of solutions are paired as anchor and positive, while nodes absent from all solutions are designated as negatives.
Training Instances: We do not expect this task to generalize between problem classes and/or sets of varying complexities of a given problem class. For instance, there is no reason for an LP gap of 70% to be the same between easy vs. hard SC instances. Figure 4(B) shows the distribution of integrality gap across different problem types and difficulties. Note that the integrality gap can be anywhere from 5% to 95% with wide distributions in between. As such, there is no magic constant that one could use at all times heuristically, which makes integrality gap prediction a deliberate learning task. For training, we consider CA (very-easy, easy, medium), SC (easy, medium, hard), and GIS (easy, medium, hard) with 50 instances for each. In total, we obtain a total of 450 training instances. Note that this is a considerably smaller training set than initially used for training Forge embeddings.
Label Collection: To generate training labels, each training instance is solved using Gurobi with a time limit of 120 seconds. Note that this does not require solving the instance to optimality which can be computationally challenging. The numeric label is defined as the ratio (or gap) between the integer program solution at the time-out value and the LP relaxation solution. As mentioned, overestimating the integrality gap (or underestimating for maximization problems) can lead to suboptimal solutions, as the solver may terminate prematurely. In contrast, underestimating the gap is generally acceptable, as it still facilitates faster solve times without compromising solution quality. This asymmetry is accounted for by adopting a conservative labeling strategy by setting a 120-second timeout when collecting labels. This often results in an underestimation of the true integrality gap, particularly for more challenging instances. By doing so, we reduce the risk of the model overestimating the gap, thereby improving the reliability of the predicted cuts.
Supervised Fine-Tuning: Given the labeled data, a dense prediction head is added to the pre-trained Forge model that takes as input the codewords assigned to each node and outputs a real number (Figure 4(A)). The mean readout across all nodes is used as the predicted integrality gap. As a regression task, this is trained with the mean absolute error loss in an end-to-end manner.
5.2.1 Numerical Results: Integrality Gap Prediction
Test Instances & Setup: We start with the pre-trained Forge model, fine-tuned with a prediction head using the labeled data, and predict the integrality gap of 50 ‘very-hard’ instances each of CA, SC, GIS, and MVC. Given the prediction, a pseudo-cut is generated by adjusting the initial LP relaxation objective - by increasing it for minimization problems or decreasing it for maximization problems. This pseudo-cut is incorporated into the original formulation as an additional constraint, enforcing that the integral objective must exceed (or fall below) the computed pseudo-cut value.
Prediction Accuracy: We measure the deviation in terms of mean absolute error between the true integrality gap and the gap predicted by Forge. On the ‘very-hard’ instances, Forge achieves a deviation of $15.42\%$, $13.55\%$, $12.03\%$ and $19.077\%$ for CA, SC, GIS and MVC instances respectively. Note that MVC instances were not part of the train set.
While predicting integrality gaps may appear straightforward, training Forge from scratch for this task led to a $\approx 33$% increase on average in mean absolute error across all problem types compared to fine-tuning a pre-trained model. This highlights the importance of unsupervised pre-training in capturing transferable structural patterns across diverse MIP instances.
Comparison with the Commercial MIP Solver: We compare the default Gurobi solver on these ‘very-hard‘ instances with and without our predicted pseudo-cut. Figure 5 shows the primal gap averaged over these 50 ‘very-hard’ instances of each type with a time limit of 3600 seconds. Without exception, across all problem types, the use of pseudo-cuts generated by Forge embeddings consistently results in lower primal gaps over time. The gains range from 30% to 85%. The default solver attains immediate improvements early on in the search and our pseudo-cuts make these gains more pronounced immediately. Recall that this only needed 50 instances from each problem type.
Figure 5: Gurobi vs. Gurobi + Pseudo-Cuts generated by Forge. Each subplot shows the primal gap reduction averaged across 50 ‘very-hard’ instances per problem type (lower is better).
Comparison with SOTA ML Methods: To further evaluate generalization across problem types and sizes, we compare against the setup described in  Li et al. (2025), where a GNN is trained on 38,256 instances from 643 generated problem types and tested on 11,584 instances spanning 157 problem types. We use Forge as is, not using any additional training instances apart from 450 D-MIPLIB instances, and test on 17,500 previously unseen instances spanning 400 generated problem types. Forge achieves a mean deviation of 18.63% in integrality gap prediction, outperforming the reported 20.14% deviation by almost 10%.
5.3 Warm-Start Prediction & Guiding The Search
Warm Start Prediction: Recall that Forge generates embedding vectors per instance and per variable/constraint. We fine tune the variable embeddings to predict which variables are likely to be in the optimal solution. These predictions are then passed into Gurobi as variable hints. A straightforward method of predicting which variables are likely to be part of the solution is to treat it as a binary classification problem and use binary cross entropy (BCE) loss. However, this poses challenges due to the large class imbalance where most variables are not part of the solution.
To address this issue, we generate 5 feasible solutions for each instance. Variables in the instance are grouped by how many of the 5 solutions each variable was repeated in. A variable is considered a “negative” variable if and only if it appears in none of the 5 solutions. Next, we use triplet loss (Schroff et al., 2015) to fine tune the embeddings of the variables. The triplet loss is given by
$L(a,p,n)=max\{d(a_{i},p_{i})-d(a_{i},n_{i})+margin,0\}$ (5)
Here, $a$ is the ‘anchor’ node, $p$ is the ‘positive’ node, $n$ is the ‘negative’ node and $d$ is a distance function (euclidean distance in our case). Triplet loss aims to minimize the distance between the ‘anchor’ and ‘positive’ nodes while ensuring the ‘negative’ node is at least ‘margin’ distance away from the ‘anchor’ node (margin is set to 2 in our case). In our setting, as shown in Figure 4(C), all variables appearing in the same number of solutions are treated as ‘positive’ and ‘anchor’ pairs. A key challenge of triplet loss is finding good negative nodes, as picking trivially negative nodes does not aid learning. We pick variables that have not appeared in any solution but are closest to the positive node in the unsupervised embedding space as negative nodes. Finally, the Forge model is fine-tuned to predict warm start variables using a combination of triplet loss and BCE loss.
Training Instances: We collect 100 instances each from CA (easy, medium), SC (easy, medium, hard) and GIS (easy, medium) for a total of 700 training instances.
Label Collection: Each instance is solved using Gurobi and a solution pool of 5. Here, we set a larger time limit of 300 seconds since we would like to get closer to the optimal solution.
Supervised Fine Tuning: The triplet loss is used to fine tune the GNN within Forge to encourage meaningful separation in the embedding space. In parallel, a dense prediction head is added to Forge that takes as input the codewords assigned to each node (Figure 4(C)). The model is trained using both the triplet loss and binary cross entropy loss on the prediction head.
5.3.1 Numerical Results: Warm-Start Prediction
Test Instances & Setup: We start with the pre-trained Forge model, fine-tuned with a prediction head using the labeled data, and then predict warm start variables of 50 ‘medium’ instances each of Combinatorial Auction (CA), Set Cover (SC), Generalized Independent Set (GIS), and Minimum Vertex Cover (MVC). Note that MVC instances are not part of the train set.
To predict the warm start variables, we begin by generating a feasible solution using Gurobi with a 1-second timeout. The variables included in this solution are designated as seed variables. Since fine-tuning with triplet loss encourages nodes corresponding to solution variables to cluster in the embedding space, we identify additional variables whose embeddings lie within a fixed radius (set to 0.1) of the seed variables. Simultaneously, we select variables ranked in the top 5th percentile and bottom 10th percentile by the prediction head as positive and negative candidates, respectively. The intersection of variables identified by the embedding-based proximity search and the prediction head are then passed to Gurobi as solution hints.
Comparision with the Commercially MIP Solver: We compare the default Gurobi solver on these ‘medium‘ instances with and without our predicted warm start variables. Figure 6 shows primal gaps averaged over 50 instances for each of CA, SC, GIS and MVC using Gurobi versus Gurobi warm started with predicted variables with a time limit of 3600 seconds. The warm-started variants consistently achieve lower primal gaps and converge to optimal solutions significantly faster, demonstrating the practical utility of Forge in accelerating MIP solving.
Figure 6: Gurobi vs. Gurobi + Warm Starts generated by Forge. Each subplot shows the primal gap reduction averaged across 50 ‘medium’ instances per problem type (lower is better). Gurobi + Warm Starts also converges to the optimal solution significantly faster.
Figure 7: Performance comparison of three variants of PS-Gurobi: the original model, a version augmented with 64-dimensional random vectors, and a version augmented with Forge embeddings (reduced to 64 dimensions via PCA). Evaluation is conducted on 50 ‘medium’ instances each of CA and GIS problems. The Forge-augmented PS-Gurobi consistently outperforms the baseline, achieving average reductions in primal gap of 41.07% for CA and 50.51% for GISP.
Comparison with SOTA ML Methods: As an additional validation of the effectiveness of Forge embeddings, we augment PS-Gurobi (Han et al., 2023), a state-of-the-art ML-based method that already demonstrates strong performance relative to other learning-based approaches. Specifically, we augment the input node features for variables and constraints in PS-Gurobi by concatenating the unsupervised embeddings generated by the pre-trained Forge model.
Since Forge embeddings for each node are 1024-dimensional, we apply Principal Component Analysis (PCA) to reduce the dimensionality to 64 dimensions, thereby minimizing architectural modifications to the original model. To further ensure that any observed performance gains are attributed to the semantic content of the Forge embeddings rather than increased model capacity, we also conduct a control experiment by augmenting PS-Gurobi with 64-dimensional random vectors. This comparison allows us to isolate the contribution of the learned embeddings to overall performance. We evaluate this setup on the same 50 instances each of the CA and GIS problems - two problem types that are common to both Forge and PS-Gurobi. Results are shown in Figure 7 where PS-Gurobi + Forge consistently outperforms the standalone PS-Gurobi with the gain in primal gaps being 41.07% and 50.51% for CA and GIS respectively. Interestingly, PS-Gurobi with random vectors occasionally surpasses the original model, likely due to the significant increase in input dimensionality - from 4 and 6 features for constraints and variables, respectively, to 68 and 70 - resulting in a larger model capacity that can yield marginal performance gains.
6 Related Work
Recent advances in applying machine learning to combinatorial optimization, particularly Mixed Integer Linear Programming (MILP or MIP), have led to a diverse set of approaches. These methods can be broadly categorized based on their focus areas, such as solver configuration, branching strategies, heuristic design, and generalization across problem types.
Learning-Based Enhancements to MILP Solvers Several works have explored integrating supervised and reinforcement learning into traditional MILP solving pipelines. The survey by Zhang et al. (2023) categorizes these methods into two main groups: those enhancing the Branch-and-Bound process (e.g., branching variable prediction, cutting plane selection, node selection), and those improving heuristic algorithms like Large Neighborhood Search (LNS), Feasibility Pump, and Predict-and-Pick. These models are typically trained end-to-end, with reinforcement learning often relying on imitation learning.
In the context of solver configuration, Hosny & Reda (2024) propose predicting solver parameters by leveraging similarities between problem instances. Their key assumption is that instances with similar costs under one configuration will behave similarly under others. Features include pre-solve statistics and tree-based metrics, with a triplet loss guiding the learning process using solved instance objectives. Kadioglu et al. (2010) propose the ISAC framework that takes a clustering-based approach to instance-specific algorithm configuration. It uses G-means to cluster problem instances and assigns configurations based on cluster membership. The method considers domain-specific features like cost-density ratios and root cost metrics for problems like SCP, MIP, and SAT.
To improve generalization, Boisvert et al. (2024) propose a generic representation for combinatorial problems using abstract syntax trees with 5 node types - variables, constraints, values, operators and a model node. While expressive, this approach results in large, computationally expensive graphs.
Multi-Task and Generalist Models Efforts to unify learning across tasks and problem types have also emerged. Cai et al. (2025a) introduce a multi-task representation learning framework for MIP, training a shared backbone across tasks such as backdoor prediction and solver configuration prediction, followed by fine-tuning for specific problem types. Their method uses a bipartite graph representation, Graph Attention Networks (GAT), and contrastive loss, and is evaluated on problems like CA, MVC and Maximum Independent Set (MIS). Similarly, Drakulic et al. (2024) present GOAL, a generalist agent for combinatorial optimization. It avoids GNNs, instead using mixed attention over edge and node matrices derived from bipartite graphs. Li et al. (2025) propose an LLM-based evolutionary framework that can generate a large set of diverse MILP classes and can be fine tuned to predict integrality gaps and branching nodes.
Graph Neural Networks for Branching and Heuristics Graph-based representations have become standard for encoding MIP instances. One of the earliest work by Gasse et al. (2019) uses GNNs to learn strong branching policies, introducing a bipartite graph structure and dual half-convolutions for message passing between constraints and variables. Chen et al. (2024) revisit GNN for MIPs and show that higher-order GNNs can overcome limitations identified via the 1-Weisfeiler-Lehman test, making all instances tractable for message passing. Cantürk et al. (2024) introduce improvements to the standard GNN workflow for CO so that they generalize on instances of a larger scale than those used in training and propose a two-stage primal heuristic strategy based on uncertainty quantification to automatically configure how solution search relies on the predicted decision values.
Along the lines of Backdoor learning, Cai et al. (2024) use Monte Carlo Tree Search to identify effective backdoors, training a GAT to score variables. Ferber et al. (2022) propose pseudo-backdoors, using one model that characterizes if a subset of variables is a good backdoor and another model to predict whether prioritizing this subset would lead to a smaller run time.
Learning Heuristics and Large Neighborhood Search (LNS) Policies A growing body of work focuses on learning heuristics, particularly for LNS. Huang et al. (2023) use expert hueuristics to create training data followed by random perturbations to create ‘negative’ samples. Then contrastive learning is used to train GATs to predict node probabilities. Other works by Wu et al. (2021) and Song et al. (2020) use deep reinforcement learning to learn destroy operators or decompositions, with rewards based on objective improvements. Khalil et al. (2017) model the success of heuristics at specific nodes by examining instance based characteristics and use logistic regression over a rich feature set, including LP relaxation and scoring metrics. Concretely, they learn a binary classifier to answer whether some heuristic H will find an incumbent solution at a given node N.
Problem Specific Solutions The vehicle routing problem (VRP) has garnered special attention from the community. Zhou et al. (2023) introduce a meta-learning framework for VRPs, enabling generalization across problem sizes and distributions. Berto et al. (2024) explore ML solutions for different kinds of VRPs like those including backhauls, multi-depots, duration limits, mixed backhaul, line hauls etc.. They use a common encoder for all VRP types with global attributes for problem type and local node attributes to capture customer specific attributes like location and demands.
Another problem that has garnered attention is the constraint satisfaction problem. Tönshoff et al. (2023) use GNNs to predict soft assignments, with reinforcement learning rewards based on constraint satisfaction improvements. Duan et al. (2022) propose a contrastive learning framework that generates label-preserving augmentations for SAT problems. These include techniques like unit clause propagation, pure literal elimination, and clause resolution, ensuring that the satisfiability of the instance remains unchanged while enhancing the model’s robustness. Shafi et al. (2025) introduce Graph-SCP, a method that leverages features extracted from both bipartite and hypergraph representations of SCP instances. A GNN is then trained with these features to predict a promising subproblem where the optimal solution is likely to reside. This predicted subproblem is passed to a solver, effectively accelerating the overall solution process.
Unsupervised Approaches Unsupervised learning has also been explored in various forms. Karalias & Loukas (2020) introduce a framework that learns a probability distribution over nodes, optimizing a loss that bounds the probability of finding a solution. These are then decoded using a derandomization process. Bu et al. (2024) build on the work by Karalias & Loukas (2020) by formalizing objective construction and derandomization strategies. They derive explicit formulations tailored to a range of combinatorial problems, including facility location, maximum coverage, and robust graph coloring. Sanokowski et al. (2024) provide an approach for solving combinatorial optimization problems without labeled data by leveraging diffusion models to sample from complex discrete distributions. Their method avoids the need for exact likelihoods by optimizing a loss that upper bounds the reverse KL divergence. While Forge falls in the domain of unsupervised approaches, we differ in that our goal is to not solve a given instance in an unsupervised manner, but rather to learn the graphical structure of various MIP instances in an unsupervised manner followed by supervised fine-tuning to aid in finding the solution.
7 Conclusion & Future Work
We introduce Forge, an unsupervised framework for learning structural representations of MIP instances without requiring access to their solutions. By using a vector-quantized graph autoencoder, Forge effectively captures both local and global structural properties - an essential capability for addressing combinatorial optimization (CO) problems. The discrete codes produced by the vector quantization process behaves as a vocabulary of MIP instances. We demonstrate that these codes can meaningfully cluster previously unseen MIP instances and can be fine-tuned for diverse downstream tasks, including pseudo-cut generation via integrality gap prediction and warm-start variable prediction. Remarkably, a single model with only 3.25 million parameters is used across both tasks and multiple problem types and difficulty levels. We further show that the embeddings produced by Forge can be seamlessly integrated into existing machine learning pipelines for CO, yielding measurable improvements in solver performance. Looking ahead, the instance-level embeddings generated by Forge offer promising opportunities for reinforcement learning-based approaches, such as guiding subproblem selection in Large Neighborhood Search frameworks. Exploring such integrations represents a compelling direction for future work. To support reproducibility and encourage further research, we release all code and pre-trained models.